# Extended Fractional Chern Insulators Near Half Flux in Twisted Bilayer Graphene Above the Magic Angle

### v2

@@ V Discussion of possible mechanisms for extended FCI
We turn now to the $8/3$ fractional state which surprisingly persists over a larger range of density than any integer state. A $18.5$ T linecut near 2/5 flux (Fig. 2B, C) traverses the three largest pockets of quantization in the Landau fan diagram, centered at $n/n_{s}=-1.3,1.9$, and $2.6$. These occur upon doping into the next band above the Hofstadter gaps $(0,-4)$, $(0,+4)$, and $(+4,-4)$, respectively. Assuming the total Chern number ±$\pm 8/3$ extracted from measured Hall resistance is the sum of the Chern number $\Delta t$ of the FCI and that of the gapped integer state below, the Chern numbers of the three FCI states would be $4/3=-8/3-(-4)$, $-4/3=8/3-4$, and again $4/3=-8/3-(-4)$, respectively. In Fig. 3A, we mark with labels $\alpha$, $\beta$, and $\gamma$, respectively, the parent spin-polarized Hofstadter subbands at 2/5 flux within our single-particle model that we believe host the FCI states. If there is minimal mixing of bands from interactions, we are in each case doping into a spin-polarized (all blue in the figure) but valley-degenerate band. Thus the 4/3 Chern number referenced above could represent two copies of a $|\Delta t|=2/3$ FCI. It could instead be one copy of 4/3, or even four copies of 1/3 if we could recover a larger degeneracy.
@@
Specifically, subband $\alpha$ and its corresponding Hall plateau are enigmatic. We observe an enormous region of quantization ($\Delta n/n_{s}\sim 0.5$), larger than any other plateau—integer or fractional—between 00 and $28$ T (see Fig. S8). An incompressible gap is expected to follow a well-defined Středa relation with slope equal to the total Chern number of the occupied bands. However, the breadth of this plateau stymies assigning a specific ($s$, $t$). To retain quantization as carrier density is tuned, the excess carriers must go into localized states. What states are doping into here? The obvious culprit would be localized states from disorder. While we cannot explicitly rule out disorder, it is unlikely that localized states from disorder would specifically extend these $\pm$8/3 fractional gaps so much more dramatically than other gaps in the spectrum. In Sec. S12 and Sec. S13, we discuss the effects of domain walls and spin-dependent transport, respectively. We consider both scenarios possible but unlikely as explanations of our data.
@@
This mechanism has been invoked to explain reentrant quantum Hall in 2D electron gases [38, 39], but for the present system we have two complicating factors. First, the presence of a moiré potential might be expected to favor forming a Wigner crystal not at a continuous range of density but specifically at rational fractional fillings of the moiré. We instead observe sustained quantization over a range of density that would encompass many simple fractions. Second, this plateau occurs in a regime of two valley-degenerate Hofstadter subbands, with bandwidth comparable to Zeeman splitting. There could thus be more than one way for the multiple flavors to combine into a fractional state, reminiscent of $\nu=2/3$ fractional quantum Hall in a seminconductor double quantum well [49]. Understanding the nature of the extended FCIs and the conditions for a subband to fractionalize warrant further experimental and theoretical investigation. If the electronic solid picture we suggest is correct, disentangling the precise nature of the solid, including whether it is partially or entirely composed of a second flavor of the composite fermions themselves, will be an important challenge.
@@
Acknowledgments We would like to thank Steve Kivelson, Julian May-Mann, Trithep Devakul, Yves Kwan, Tomohiro Soejima, Patrick Ledwith, Ben Feldman, Matt Yankowitz, Dave Cobden, Xiaodong Xu, Wei Pan, Jonah Herzog-Arbeitman, Nisarga Paul, and Sayak Bhattacharjee for fruitful discussions. Sample preparation, measurements, and analysis were supported by the US Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division, under Contract DE-AC02-76SF00515. Device fabrication was performed at the Stanford Nano Shared Facilities which is supported by the NSF under award ECCS-2026822. D.G.-G. acknowledges support for supplies from the Ross M. Brown Family Foundation and from the Gordon and Betty Moore Foundation’s Emergent Phenomena in Quantum Systems (EPiQS) Initiative through grant GBMF9460. X.W. acknowledges financial support from National MagLab through Dirac fellowship, which is funded by the National Science Foundation (Grant No. DMR-1644779) and the state of Florida. O.V. was funded in part by the Gordon and Betty Moore Foundation’s EPiQS Initiative Grant GBMF11070 and acknowledges support from the National High Magnetic Field Laboratory funded by the National Science Foundation (Grant No. DMR-2128556) and the State of Florida. K.W. and T.T. acknowledge support from JSPS KAKENHI (Grant Numbers 19H05790, 20H00354 and 21H05233). Part of this work was performed at the Stanford Nano Shared Facilities (SNSF), supported by the National Science Foundation under award ECCS-2026822.
@@
S1 Generalized Diophantine and Definitions
@@
Both an incompressible (fractional) quantum Hall liquid and a Hall crystal are described by the case where $\nu_{s}\neq 0$ and $\eta_{s}=0$. In the latter case, the system spontaneously breaks translation symmetry with density fluctuations set by $\rho_{\phi}$ while maintaining a non-zero Hall conductance. In an anomalous Hall crystal, time reversal symmetry breaks spontaneously in addition to translation symmetry breaking. A Wigner crystal, however, is described by $\nu_{s}=0$ and $\eta_{s}\neq 0$ and therefore must have vanishing Hall conductance.
@@
A partial Hall crystal is, in some sense, an intermediate phase combining the properties of both. It is described by $\nu_{s}\neq 0$ and $\eta_{s}\neq 0$. Therefore, it will have a non-zero Hall conductance and spatially modulated electron density. Given the Wigner crystal background, a partial Hall crystal will be compressible if the spatial modulation of the electron density is not strongly pinned to a disorder potential.
S2 Transport measurements
We measured the sample in the SCM-32T system at the National High Magnetic Field Laboratory (MagLab). Although the hybrid magnet had a nominal maximum field of 32 T, because we were one of the early users of the system, we were restricted to a maximum field of 28 T. We used the top-loading dil fridge insert with a base temperature of 50 mK at the mixing chamber plate. The sample probe had 16 DC measurement wires, so we were not able to measure every contact pair in our device.
The probe did not feature cold low-pass filters. We included low-pass filters (borrowed from another user and described in Ref. [50]) at the breakout box, however there was still RF pickup between the breakout box and the sample. This was most evident at 8 PM every day, when a nearby radio station turned down their output power and our measurement noise decreased. For example, in Fig. 1, we crossed 16.8 T at 8 PM, where there is a noticable increase in sharpness of features.
In addition to RF noise blurring our measurements, they also suffered from low-frequency noise within the measurement instruments themselves. This noise appeared in the current measurement for Fig. 2 (Fig. S6), but not in the associated voltage measurements for all densities/fields, meaning that the noise was not in the current itself. Thus, for Fig. 2, we do not use the current measurement, but instead divide by a constant 1 nA.
[…]
We took the first set of measurements on many contact pairs at once (Figs. S2 and S3). The Hall measurements in Fig. 2 were taken in this set. We then found that only measuring one contact pair at a time greatly reduced noise. We therefore took the measurement in Fig. 1, consisting only of a single longitudinal resistance measurement, separately.
We show a comparison between the data taken at MagLab with data taken in our own system at Stanford in Fig. S4. The system at Stanford both has a lower base temperature and high quality filters at the mixing chamber stage. There are two stages of filtering: the wires are first passed through a cured mixture of epoxy and bronze powder to filter GHz frequencies, then low-pass RC filters mounted on sapphire plates filter MHz frequencies. As a result, the Landau fan appears much sharper. The resistivity near charge neutrality is nearly an order of magnitude higher in our measurements at Stanford compared to those at MagLab.
There is one notable difference between how the measurements were performed at Stanford vs MagLab. At Stanford, our contacts became highly resistive when the graphite back gate was held at ground, so we fixed the graphite back gate to 1.5 V for the measurements. At MagLab, we found that the contacts were an order of magnitude lower resistance than what we saw at Stanford, so we left the graphite back gate at ground. The difference in contact resistances was likely a function of the higher effective temperature.
S3 Cluster filtering
The unfiltered data from Fig. 2 can be seen in Fig. S21. To filter the data, we begin by identifying points where the Hall resistance $R_{xy}$ falls within some window of our desired value (±$\pm 1\%$ of ±$\pm 3h/8e^{2}$ in this case). Binary dilation is applied to the windowed data such that nearest-neighbors within this masked array are now included. Each contiguous cluster of the dilated data set is then identified, and, if it does not contain above a threshold number of points, the original undilated points are omitted from the final filtered data set. Because of the dilation and generous threshold, this filtering primarly serves to remove isolated data points with no next-nearest-neighbor point within the window.
S4 Brown-Zak oscillations
@@
At rational values of $\Phi/\Phi_{0}=p/q$, the system can be described in terms of Bloch states in a $q$-times enlarged magnetic Brillouin zone [51, 52]. In transport, this manifests as oscillations in conductivity at simple field fractions and constant density. These “Brown-Zak” oscillations have been observed in graphene/hBN superlattices[53, 54, 55]. The homogeneity and amount of disorder in the sample affects the maximum $q$ for which oscillations will be visible, because as $q$ increases, the size of the magnetic Brillouin zone increases relative to the carrier mean free path [54].
@@
We observe Brown-Zak oscillations at $\Phi/\Phi_{0}=1/3$, $2/5$, $1/2$, and $3/5$ (panel B, horizontal dotted lines). Of note, the Brown-Zak oscillations at $1/2$ flux appear only between $n/n_{s}=\pm 2$. The density-dependent amplitude of the oscillations is a probe of the carrier group velocity and therefore the miniband width [54]. So, where we do not see Brown-Zak oscillations, we should expect flatter bands and thus stronger electron correlations. In general, we observe more and stronger Brown-Zak oscillations on the electron side of charge neutrality compared to the hole side, consistent with the basic principle that the hole bands are flatter than electron bands in TBG.
@@
We observe a similar set of Brown-Zak oscillations in contact pairs 16-17, 7-8, and 17-18 (Figs. 1 and S2). Contact pair 6-7 has generally blurrier features, and has broad oscillations only at $\Phi/\Phi_{0}=1/2$ and $1/3$. Contact pair 13-14 does not show clear oscillations, likely both because of inhomogeneity and also closer proximity to the magic angle.
It would be convenient if the Brown-Zak oscillations were ideally sharp features in field, because their broadening would be a function of twist angle inhomogeneity. Then, assuming Gaussian disorder (a likely wrong assumption), it would be straightforward to compute $\sigma_{\theta}$ from the FWHM of the BZ oscillations via
$\sigma_{B}=\frac{4h\theta}{\sqrt{3}ea^{2}}\left(\Phi/\Phi_{0}\right)\sigma_{\theta}.$
Unfortunately, in Fig. 1, the width of the oscillations in the oscillations is not set by twist angle inhomogeneity. This can be readily ascertained from the fact that the features at $\Phi/\Phi_{0}=3/5$ are significantly sharper than those at $1/2$.
Near $\Phi/\Phi_{0}=p/q$, the Bloch electrons feel an effective field [53]
$B_{\mathrm{eff}}=B-\frac{\Phi_{0}}{A}\frac{p}{q}.$
This is an aspect of Hofstadter’s butterfly’s fractal nature.
S5 Single-particle bends, SBCI, Hofstadter ferrogmagnetism
As demonstrated in Ref. [36], not all relatively straight lines in field/density with low resistance are Str̆eda gaps. Instead, relatively straight-line features of low resistivity can appear where one anisotropy-smeared butterfly spectrum is coincident with another. These false Str̆eda lines have a characteristic behavior at intersections with lines from the other doubled butterfly. Many of the features within our measurements follow the same pattern (see Fig. 3), and we attempt to only place lines in Fig. 1B where we are confident that there is a gap.
We observe a few Str̆eda lines with fractional $s=\pm 1/2$ and $t=\pm 3$ near half flux (red dotted lines in panel B). As they appear to emit from $n/n_{s}=\pm 2$ at half flux, we would describe them as (±$(\pm 2,\pm 3)_{1/2}$. Such lines have been described as symmetry-broken Chern insulators in the literature [19], and we observe at least one of them in all of the contact pairs that we measured. Curiously, in another contact pair, a symmetry-broken Str̆eda line $(13/3,-4)=(3,-4)_{1/3}=(7/3,-4)_{1/2}$ runs between one third and one half flux.
We observe states with $s+t=\pm 4$ projecting from integer $s$ (green dashed lines in Figs. 1 and S2). In all contact pairs, they tend to extend to lower field on the hole side compared to the electron side (Fig. S2). In samples from the literature near the magic angle, they extend to zero or near zero field [3, 56, 47, 22, 57, 7, 8, 9]. These lines are described theoretically as correlated Hofstadter ferromagnetic states in Ref. [42], meaning they have polarized spin and valley degrees of freedom of interaction-modified bands.
S6 Str̆eda line fit procedure
@@
The raw data for Figs. 1, S2, and S3 are in gate-voltage vs magnetic field. We transform the data into $n/n_{s}$ vs $\Phi/\Phi_{0}$ by placing points along sharp lines with known $(s,t)$ by-eye. Brown-Zak oscillations additionally constrain $\Phi/\Phi_{0}$. We then use least squares to fit parameters $m_{g}$, $b_{g}$, and $B_{0}$, where $n/n_{s}=m_{g}V_{g}+b_{g}$ and $\Phi/\Phi_{0}=1$ at $B=B_{0}$. The resulting fit parameters are shown in Table S1 and schematically with our device geometry in Fig.S5. Quantum capacitance, owing to its dependence on density of states, precludes an accurate global conversion between gate voltage and density. The fits should be assumed to be at least a few percent wrong where the density of states is low.
@@
After fitting the global conversion parameters to lines with known $(s,t)$, we were able to assign $(s,t)$ to the remaining straight-line features. Most of the lines could be easily assigned by-eye, and the results for all contact pairs are shown in Fig. S2. We note that not all relatively straight-line minima in $R_{xx}$ correspond to quantum Hall Str̆eda lines, as noted in Ref. [36]. We attempted to correctly label only Str̆eda lines in Figs. 1 and S2.
Because of the $8/3$ FCI states nearby, we expected to find Str̆eda lines with slope $8/3$ near half flux. As shown in Fig. S9, the candidate lines are actually better fit by $s=\pm 1/2$, $t=\pm 3$. These lines appear closely associated with half flux, so are better described by (±$(\pm 2,\pm 3)_{1/2}$.
@@
Table S1: Voltage to density conversion parameters. *: Fit to the data in Fig. 1. The rest of the rows are fits to the data in Fig. S2. It is important that $m_{g}$ and $B_{0}$ agree between the two measurements of contact pair 7 - 8, and they do. On the other hand, $b_{g}$ may vary over time as charge impurities and the like change. Areas and twist angles, along with their uncertainties, are calculated assuming zero strain. We know there is uniaxial heterostrain of order $0.2$% in this sample, so the uncertainty in $\theta$ is underestimated.
S7 Other contact pairs
Fig. S5 shows a schematic of the device with measured twist angles overlaid. We can make a rough estimate of which regions in the device have inhomogeneity/different twist angles based on the measured twist angles along with the sharpness of features. For instance, contact pairs 6 - 16 and 6 - 7 have both a higher measured twist angle but also show very blurry features in transport. Hence, we draw the blue region. Contact pairs 17 - 18 is slightly blurry, and has a slightly higher twist angle than its neighbors, but contact pair 8 - 18 is very blurry and has a significantly higher twist angle, hence the orange region. These regions are only our guesses.
@@
S8 Detail of Landau level reset
Figs. S10, S11, and S12 magnify the bending Landau minifan near $n/n_{s}=2$, $\Phi/\Phi_{0}=1/2$ from Figs. 1, S2, and S3.
In Fig. S10, red dotted lines mark bending, low-resistance states. The green dashed $(1,3)$ line is not visible as a minimum in $\rho_{xx}$. Instead, we include it to highlight that there is a kink in the red dotted lines where it would intersect them. This line corresponds to an integer filling $\nu=1$ of the band at half flux. After the reset at $n/n_{s}=3$ (half filling of the band, or $\nu=2$), we observe Str̆eda lines $(3,\pm 6)_{1/2}$ and $(3,\pm 10)_{1/2}$.

### v3

@@ V Discussion of possible mechanisms for extended FCI
We turn now to the $8/3$ fractional state which surprisingly persists over a larger range of density than any integer state. A linecut at $18.5$ T, near 2/5 flux (Fig. 2B, C), traverses the three largest pockets of quantization in the Landau fan diagram, centered at $n/n_{s}=-1.3,1.9$, and $2.6$. These occur upon doping into the next band above the Hofstadter gaps $(0,-4)$, $(0,+4)$, and $(+4,-4)$, respectively. Assuming the total Chern number ±$\pm 8/3$ extracted from measured Hall resistance is the sum of the Chern number $\Delta t$ of the FCI and that of the gapped integer state below, the Chern numbers of the three FCI states would be $+4/3=-8/3-(-4)$, $-4/3=8/3-4$, and again $+4/3=-8/3-(-4)$, respectively. In Fig. 3A, within our single-particle model we denote the parent spin-polarized Hofstadter subbands at 2/5 flux that we believe host the FCI states $\alpha$, $\beta$, and $\gamma$, respectively. If interactions do not generate substantial mixing of bands, in each case we are doping into a spin-polarized (all blue in the figure) but valley-degenerate band. In this scenario, the experimentally-extracted 4/3 Chern number represents two copies of a $|\Delta t|=2/3$ FCI. Alternatively, if valley degeneracy is spontaneously broken by interactions the observed quantization could reflect a non-degenerate $|\Delta t|=4/3$ FCI. We cannot rule out the possibility that it is four copies of a 1/3 FCI, but that would require recovery of spin degeneracy in the face of a Zeeman energy comparable to subband widths.
@@
Subband $\alpha$ and its corresponding Hall plateau are particularly enigmatic. We observe an enormous region of quantization ($\Delta n/n_{s}\sim 0.5$), larger than any other plateau—integer or fractional—between 00 and $28$ T (see Fig. S15). An incompressible gap is expected to follow a well-defined Středa relation with slope equal to the total Chern number of the occupied bands. However, the breadth of this plateau stymies assigning a specific ($s$, $t$). To retain quantization as carrier density is tuned, the excess carriers must go into localized states. What states are we doping into here? The obvious culprit would be localized states from disorder. Though we cannot explicitly rule out this scenario, it is unlikely that localized states from disorder would specifically extend these $\pm$8/3 fractional gaps so much more dramatically than other gaps in the spectrum. In Sec. S11 and Sec. S13, we discuss the effects of domain walls and spin-dependent transport, respectively. We consider both scenarios possible but unlikely as explanations of our data.
@@
This mechanism has been invoked to explain reentrant (fractional) quantum Hall in 2D electron gases [46, 47, 60], but for the present system we have two complicating factors. First, the presence of a moiré potential might be expected to favor forming a Wigner crystal specifically at rational fractional fillings of the moiré. We instead observe sustained quantization over a range of density that would encompass many simple fractions. Second, this plateau occurs in a regime of two valley-degenerate Hofstadter subbands, with bandwidth comparable to Zeeman splitting. There could thus be more than one way for the multiple flavors to combine into a fractional state, reminiscent of $\nu=2/3$ fractional quantum Hall in a seminconductor double quantum well [61]. Understanding the nature of the extended FCIs and the conditions for a subband to fractionalize warrant further experimental and theoretical investigation. If the electronic solid picture we suggest is correct, disentangling the precise nature of the solid, including whether it is partially or entirely composed of a second flavor of the composite fermions themselves, will be an important challenge.
@@
Acknowledgments We would like to thank Steve Kivelson, Julian May-Mann, Trithep Devakul, Yves Kwan, Tomohiro Soejima, Patrick Ledwith, Ben Feldman, Matt Yankowitz, Dave Cobden, Xiaodong Xu, Wei Pan, Jonah Herzog-Arbeitman, Nisarga Paul, and Sayak Bhattacharjee for fruitful discussions. Sample preparation, measurements, and analysis were supported by the US Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division, under Contract DE-AC02-76SF00515. Device fabrication was performed at the Stanford Nano Shared Facilities which is supported by the NSF under award ECCS-2026822. D.G.-G. acknowledges support for supplies from the Ross M. Brown Family Foundation and from the Gordon and Betty Moore Foundation’s Emergent Phenomena in Quantum Systems (EPiQS) Initiative through grant GBMF9460. X.W. acknowledges financial support from National MagLab through Dirac fellowship, which is funded by the National Science Foundation (Grant No. DMR-2128556) and the state of Florida. O.V. was funded in part by the Gordon and Betty Moore Foundation’s EPiQS Initiative Grant GBMF11070 and acknowledges support from the National High Magnetic Field Laboratory funded by the National Science Foundation (Grant No. DMR-2128556) and the State of Florida. K.W. and T.T. acknowledge support from JSPS KAKENHI (Grant Numbers 19H05790, 20H00354 and 21H05233). Part of this work was performed at the Stanford Nano Shared Facilities (SNSF), supported by the National Science Foundation under award ECCS-2026822. A portion of this work was performed at the National High Magnetic Field Laboratory, which is supported by the National Science Foundation Cooperative Agreement No. DMR-1644779 and the State of Florida.
Author Contributions J.F. and A.S. devised the project, performed data analysis, and wrote the manuscript with input from all authors. J.F. and L.K.R. fabricated the devices and performed measurements. J.K. X.W. and O.V. performed bandstructure calculations. K.W. and T.T. supplied the boron nitride crystals. M.A.K. and D.G-G. supervised the project.
Competing interests The authors declare no competing interests.
@@
S1 Transport measurements
We measured the sample in the SCM-32T system at the National High Magnetic Field Laboratory (MagLab). Although the hybrid magnet had a nominal maximum field of 32 T, as early users of the system we were restricted to a maximum field of 28 T. We used the top-loading dil fridge insert with a base temperature of 50 mK at the mixing chamber plate. The sample probe had 16 DC measurement wires, so we were not able to measure every contact pair in our device. The probe did not feature cold low-pass filters. We included low-pass filters (borrowed from another user and described in Ref. [62]) at the breakout box, however there was still RF pickup between the breakout box and the sample.
In addition to RF noise-induced blurring, our measurements also suffered from low-frequency noise within the measurement instruments themselves. This noise appeared in the current measurement for Fig. 2 (Fig. S6), but not in the associated voltage measurements for all densities/fields, meaning that the noise was not in the current itself. Thus, for Fig. 2, we do not use the current measurement, but instead divide by a constant 1 nA.
[…]
We took the first set of measurements on many contact pairs at once (Figs. S4 and S5). The Hall measurements in Fig. 2 were taken in this set. We then found that only measuring one contact pair at a time greatly reduced noise. We therefore took the measurement in Fig. 1, consisting only of a single longitudinal resistivity measurement, separately.
We show a comparison between the data taken at MagLab with data taken in our own system at Stanford in Fig. S2. The system at Stanford has a lower base temperature and high quality filters at the mixing chamber stage. There are two stages of filtering: the wires are first passed through a cured mixture of epoxy and bronze powder to filter GHz frequencies, then low-pass RC filters mounted on sapphire plates filter MHz frequencies. As a result, the Landau fan appears much sharper. The resistivity near charge neutrality is nearly an order of magnitude higher in our measurements at Stanford compared to those at MagLab.
There is one notable difference between how the measurements were performed at Stanford versus MagLab. At Stanford, our contacts became highly resistive when the graphite back gate was held at ground, so we fixed the graphite back gate to 1.5 V for the measurements. At MagLab, we found that the contacts were an order of magnitude lower resistance than what we saw at Stanford, so we left the graphite back gate at ground. The difference in contact resistances was likely a result of a difference in effective temperature between the two setups.
S2 Generalized Diophantine and Definitions
@@
A Wigner crystal ($\eta_{s}\neq 0$, $\nu_{s}=0$) will have a spatially modulated electron density but vanishing Hall conductance. An incompressible (fractional) quantum Hall liquid ($\eta_{s}=0$, $\nu_{s}\neq 0$), and a Hall crystal ($\eta_{s}=0$, $\nu_{s}\neq 0$) differ in that a quantum Hall liquid preserves translation symmetry whereas a Hall crystal spontaneously breaks translation symmetry with density fluctuations set by $\rho_{\phi}$ (an anomalous Hall crystal further spontaneously breaks time reversal symmetry). A Hall crystal differs from a Wigner crystal in having quantized Hall conductance: $\nu_{s}\neq 0$.
@@
A partial Hall crystal ($\nu_{s}\neq 0$ and $\eta_{s}\neq 0$) is an “intermediate case” combining the properties of both a Wigner crystal and a Hall crystal [48]. Therefore, it will have a non-zero Hall conductance and spatially modulated electron density. Given the Wigner crystal background, a partial Hall crystal will be compressible if the spatial modulation of the electron density is not strongly pinned to a disorder potential.
S3 Brown-Zak oscillations
@@
At rational values of $\Phi/\Phi_{0}=p/q$, the system can be described in terms of Bloch states in a $q$-times-enlarged magnetic Brillouin zone [63, 64]. In transport, this manifests as oscillations in conductivity at simple field fractions and constant density. These “Brown-Zak” oscillations have been observed in graphene/hBN superlattices[65, 66, 67]. The homogeneity and amount of disorder in the sample affects the maximum $q$ for which oscillations will be visible, because as $q$ increases, the size of the magnetic Brillouin zone increases relative to the carrier mean free path [66].
@@
We observe Brown-Zak oscillations at $\Phi/\Phi_{0}=1/3$, $2/5$, $1/2$, and $3/5$ (Fig. S3 panel B, horizontal dotted lines). Of note, the Brown-Zak oscillations at $1/2$ flux appear only between $n/n_{s}=\pm 2$. The amplitude of such oscillations depends on the carrier group velocity and therefore the miniband width [66], so absence of Brown-Zak oscillations in certain bands may indicate that those bands are flatter, with correspondingly greater influence of electron interactions. In general, we observe more and stronger Brown-Zak oscillations on the electron side of charge neutrality compared to the hole side, consistent with the familiar observation that in TBG the hole bands are flatter than electron bands.
@@
We observe a similar set of Brown-Zak oscillations in contact pairs 16-17, 7-8, and 17-18 (Fig. S4). Contact pair 6-7 has blurrier features overall, with Brown-Zak oscillations visible only at $\Phi/\Phi_{0}=1/2$ and $1/3$. Contact pair 13-14 does not show clear oscillations, likely because of both spatial inhomogeneity and stronger interactions brought by closer proximity to the magic angle.
S4 Str̆eda line fit procedure
@@
The raw data for Figs. 1, S4, and S5 were collected as a function of gate voltage and magnetic field. We transform the data into the space of $n/n_{s}$ vs $\Phi/\Phi_{0}$ by placing points along sharp lines with known $(s,t)$ by-eye. Brown-Zak oscillations additionally constrain $\Phi/\Phi_{0}$. We then use least squares to fit parameters $m_{g}$, $b_{g}$, and $B_{0}$, where $n/n_{s}=m_{g}V_{g}+b_{g}$ and $\Phi/\Phi_{0}=1$ at $B=B_{0}$. The resulting fit parameters are shown in Table S1 and schematically with our device geometry in Fig. S1. In this process we neglect the role of quantum capacitance, which depends on density of states and thus in turn on carrier density, making any global linear conversion between gate voltage and density imperfect.
@@
After fitting the global conversion parameters to lines with known $(s,t)$, we were able to assign $(s,t)$ to the remaining straight-line features. Most of the lines could be easily assigned by eye, and the results for all contact pairs are shown in Fig. S4. Not every relatively straight line in the space of density and flux along which $\rho_{xx}$ has a local minimum corresponds to a quantum Hall Str̆eda line, as already noted in Ref. [44]. In Figs. S3 and S4 we attempted to label only Str̆eda lines.
@@
Table S1: Voltage to density conversion parameters. Fits to the data, primarily from Fig. S4. The * marks a second measurement of contact pair 7 - 8 only, from Fig. 1 – we were able to decrease measurement noise by measuring only this one contact pair. It is important that $m_{g}$ and $B_{0}$ agree between the two measurements of contact pair 7 - 8, and they do. On the other hand, $b_{g}$ may vary between cooldowns as charges in the dielectrics or at interfaces migrate. In this table we report twist angles calculated assuming zero strain, despite the known presence of $\sim 0.2$% uniaxial and biaxial heterostrain in the sample. At finite twist angle, moiré cell area depends on heterostrain only to second order in strain amount, but the influence of strain is magnified by the small twist angle [45]. Neglecting uniaxial heterostrain biases our extracted twist angle by 0.001 degrees, comparable to our reported fit uncertainties. The effect of 0.2% biaxial heterostrain is substantially larger, biasing our extracted twist angle by 0.003 degrees, comparable and in some cases larger than our reported fit uncertainties. The unit cell area depends on biaxial homostrain to first order in strain amount, but without the small-angle magnification granted by heterostrain. Should biaxial homostrain be present in the sample at a level comparable to the heterostrain, neglecting it would bias our extracted twist angle by 0.001 degrees.
S5 Other contact pairs
Fig. S1 shows a schematic of the device with measured twist angles overlaid. We can roughly identify where in the device twist angle is inhomogeneous, based on the sharpness of transport features and the spatial variation of measured twist angle. For instance, contact pairs 6 - 16 and 6 - 7 have high measured twist angle together with very blurry features in transport.
Hence, we draw the blue region. Contact pairs 17 - 18 is slightly blurry, and has a slightly higher twist angle than its neighbors, but contact pair 8 - 18 is very blurry and has a significantly higher twist angle, hence the orange region. These regions are only our best estimates. We do not have any local probes to measure them directly.
@@
S6 Single-particle bends, SBCI, Hofstadter ferrogmagnetism
As demonstrated in Ref. [44], not all relatively straight lines in field/density with low resistance are Str̆eda gaps. Instead, relatively straight-line features of low resistivity can appear where one anisotropy-smeared butterfly spectrum is coincident with another. These false Str̆eda lines have a characteristic behavior at intersections with lines from the other doubled butterfly. Many of the features within our measurements follow the same pattern, and we attempt to only place lines in Fig. S3B where we are confident that there is a gap.
We observe a few Str̆eda lines with fractional $s=\pm 1/2$ and $t=\pm 3$ near half flux (Fig. S3 red dotted lines in panel B and right panels of Fig. S4). As they appear to emit from $n/n_{s}=\pm 2$ at half flux, we would describe them as (±$(\pm 2,\pm 3)_{1/2}$. They do not have a slope of $8/3$ (Fig. S7). Such lines have been described as symmetry-broken Chern insulators in the literature [19], and we observe at least one of them in all of the contact pairs that we measured. Curiously, in contact pair 16 - 17, a symmetry-broken Str̆eda line $(13/3,-4)=(3,-4)_{1/3}=(7/3,-4)_{1/2}$ runs between one third and one half flux.
We observe states with $s+t=\pm 4$ projecting from integer $s$ (green dashed lines in Figs. S3 and S4). In all contact pairs, they tend to extend to lower field on the hole side compared to the electron side (Fig. S4). In samples from the literature near the magic angle, they extend to zero or near zero field [3, 58, 56, 22, 59, 7, 8, 9]. These lines are described theoretically as correlated Hofstadter ferromagnetic states in Ref. [50], meaning they have polarized spin and valley degrees of freedom of interaction-modified bands.
S7 Detail of Landau level reset
Figs. S8, S9, and S10 magnify the bending Landau minifan near $n/n_{s}=2$, $\Phi/\Phi_{0}=1/2$ from Figs. 1, S4, and S5.
In Fig. S8, red dotted lines mark bending, low-resistance states. The green dashed $(1,3)$ line is not visible as a minimum in $\rho_{xx}$. Instead, we include it to highlight that there is a kink in the red dotted lines where it would intersect them. This line corresponds to an integer filling of the band at half flux. After the reset at $n/n_{s}=3$ (half filling of the band), we observe Str̆eda lines $(3,\pm 6)_{1/2}$ and $(3,\pm 10)_{1/2}$.